OpenStudy (anonymous):

distance between (0,0) and (8,2)

OpenStudy (anonymous):

=sqrt(64+4)=sqrt(68)=2sqrt(17)

OpenStudy (anonymous):

m=1/4

OpenStudy (anonymous):

\[2\sqrt{17}\]

OpenStudy (anonymous):

there hoe bag:)